# Towards Human-AI Synergy in UI Design: Enhancing Multi-Agent Based UI Generation with Intent Clarification and Alignment

### v1

@@ 3.1. Findings
Our formative interviews confirmed the limitations of current input and output modalities used by designers, highlighting the need to address design aspects that better support the generation process. Unlike commercial tools, our PrototypeAgent focuses on implicit design intent clarification, intent-design alignment, and ensuring coherence between design elements throughout the prototype generation process. Table 1 in Section 2.1 compares popular design and generative tools used in participant workflows, helping to illustrate our findings and the advantages of our proposed method. We have identified five key findings as follows:
F1. Seamless and Flexible Integration of Knowledge Bases While there are numerous design search tools and generative models available, a design space related to the use of knowledge bases still needs to be considered. Participants in our study highlighted that knowledge bases should be embedded directly into the design workflow, rather than existing as isolated resources. This integration can enhance the accuracy and automation of content generation and use, while also reducing the risks of generating irrelevant or disconnected designs. As one participant remarked, “I regularly discover designs or elements that match what I’m looking for. However, seamlessly integrating them into my design process remains difficult. AI-powered tools should make this easier.”
In addition to seamless integration, participants highlighted the need for systems that support the ongoing expansion of their knowledge bases. They expressed a desire for scalability, allowing designers to build upon and customize their existing resources as their needs evolve. To address this, our approach incorporates a dynamic, non-training method for utilizing the knowledge base: retrieval-augmented generation (RAG). This ensures continuous augmentation of the designer’s knowledge base without requiring extensive retraining, making the system both flexible and adaptive.
F2. Design Interaction Modality Requirements Our findings indicate that design interaction modalities should prioritize controllability in terms of functionality and layout, as well as the editability of outputs, as outlined in Table 1 Input and Output.
@@
Traditional design tools often require manual creation of designs, typically in SVG format. Emerging generative tools offer designers the ability to input natural language descriptions and high or mid-fidelity design images (e.g., screenshots). However, interviews with participants revealed that the ability to precisely and controllably describe their design intent is a crucial requirement for effective input. As one designer expressed, “Text-only or image-only input doesn’t fully capture what I’m imagining, especially regarding the functionality of each component, which limits their practical use.” Participants noted the advantages of combining wireframe screenshots with text descriptions, which not only provide control over functionality and layout but also preserve creativity and the automation benefits of AI-powered generation tools.
@@
Regarding Output, participants emphasized that editability is essential. While SVG format allows for easy modifications, other formats, such as code (e.g., Vercel’s V0), are challenging for designers lacking programming skills. As one participant shared, “Getting a design from MidJourney is amazing, but the real work begins with the day of editing that follows.” This underscores the need for output formats that support easy editing and refinement.
F3. Bridging AI and Designer Intent through Implicit Intent Clarification Designers expressed that AI and designers should collaborate to enhance each other’s work, especially when it comes to clarifying design intent. One of the key areas highlighted is the need for AI tools to assist in refining the initial input provided by designers. Due to unfamiliarity with prompt structures or unclear design objectives, designers often start with simple inputs like “design a landing page” without specifying the granular details of their requirements. AI should work with designers to clarify and refine these inputs.
@@
As one participant noted, “I believe AI understands the kind of instructions it needs better than I do. If it could provide feedback, I could express my design goals more quickly and more accurately.” This emphasizes the importance of automated intent clarification as part of the AI-human collaboration process, aligning with the concept of achieving common ground in communication theory (Clark, 1991).
To address this, our system incorporates an Implicit Intent Clarification step, which refines prompts by regenerate them into specific design aspects. As highlighted in Table 1 under Implicit Intent Clarification, this is an important aspect of AI assistance that is currently missing in other AI generation tools.
F4. Decoupling the Generation Process for Human-Centered Intent-Design Alignment Studies (Dang et al., 2023) have highlighted the lack of transparency and explainability in end-to-end generation methods, particularly in achieving intent-design alignment from themes to local components. Designers often need to make adjustments—changing design styles, editing text, or replacing images—but the opacity of these methods forces them to “guess” the correct input phase, wasting time and reducing accuracy.
@@
As shown in Table 1 under Human-Centered Intent-Design Alignment, current tools lack clear explanations of their generation steps. Designers called for more human-centered mechanisms to better understand and refine the process. One participant remarked, “If AI tools could reveal their thought processes and allow editing, it would greatly simplify tailoring the results to our needs.”
Our approach addresses this by decoupling the generation process and integrating intent-design alignment. This allows designers and AI to collaboratively refine intermediate guidance and results. Using natural language as DSL for global and local descriptions, our method promotes clarity, reduces ambiguity, and supports both large-scale and component-level alignment.
F5. Ensuring Design Coherence Through Generation Processes Maintaining consistency in design elements—such as color, style, and aesthetics—across a project is often a repetitive and labor-intensive task. As shown in Table 1 under Design Coherence Among Elements (Semantic + Style), traditional design tools require manual intervention to achieve visual cohesion. While newer generative tools can use text descriptions or reference images to assist with this, their ability to maintain consistency remains unclear. As one participant noted, “Creating a unified look for a new app involves a lot of repetitive work.” This underscores the need for tools that can automatically maintain design consistency.
Our approach addresses this challenge by employing an agent as a central controller that manages design consistency across various components. It not only ensures coherence but also explicitly presents the natural language content describing design consistency, allowing designers to interact with and refine the results efficiently.
@@ 4.1. Overview
To improve the findings identified in Section 3 and actualize the design space outlined, we introduce PrototypeAgent: a system that harnesses multi-agent collaboration for the explainable generation of UI prototypes. This system aims to enhance mutual disambiguation in Human-AI Collaboration for creating high-fidelity prototypes by offering a decoupled generation process.
While large language models excel in contextual understanding and response generation, UI generation poses unique challenges that require deep domain knowledge to ensure high-quality design. A simple end-to-end approach may fall short of capturing the complexities of UI design intent. Therefore, we adopt a divide-and-conquer, top-down strategy that decouples the generation process into sub-steps. A central managing agent oversees these steps to ensure coherence across the design. By using natural language descriptions as a DSL, our method enables designers and AI to collaboratively understand and refine intermediate results.
As depicted in Fig. 2, PrototypeAgent integrates knowledge bases with four specialized agents: Theme Design Agent $A_{theme}$, Text Content Agent $A_{text}$, Image Content Agent $A_{img}$, and Icon Agent $A_{icon}$. We first collected two knowledge bases offline to import relevant design knowledge, corresponding to F1. Seamless and Flexible Integration of Knowledge Bases in the Section 3.1. The online system commences its operation by processing design inputs, including a prompt and UI layout, as specified in F2. Design Interaction Modality Requirements. The Theme Design Agent initiates the process by generating a high-level theme description for F3. Implicit Intent Clarification and corresponding image, setting the stage for a consistent design narrative in line with F5. Ensuring Design Coherence.
Functioning as the primary director, $A_{theme}$ delegates specialized tasks to $A_{text}$, $A_{img}$, and $A_{icon}$ for the detailed creation of each design component, addressing F4. Human-Centered Intent-Design Alignment and F5 respectively. The effective coordination between the central and supporting agents, facilitated by a cache pool, is crucial for maintaining the design’s contextual integrity.
This methodology fosters clarity, disambiguation and iterative interactions, allowing PrototypeAgent to progressively refine the design. Designers are thus provided with an environment that not only generates designs but also supports iterative refinements to align closely with their intent.
@@ 4.2. Knowledge Base Construction
Employing domain-specific knowledge can harness the creativity of Large Language Models (LLMs) while enhancing the quality of generation (Wang et al., 2023b; Lu et al., 2024). We collected two kinds of knowledge, namely UI knowledge (pairs of layout/theme descriptions with local component descriptions) and Icon Knowledge (Pairs of icon SVG code with semantic descriptions) for $A_{theme}$ and $A_{icon}$.
@@ 4.2.1. UI Knowledge Base
1) UI Composition Knowledge. Rico Dataset (Deka et al., 2017) is one of the most comprehensive open-source UI datasets available, which contains around 66k distinct UIs from over 9.7k Android apps across 27 categories. This dataset includes the UI screenshots and their view hierarchy information, which expose UI elements used, their attributes like text, bounds and class, and the composition of these UI elements. We extracted class and bounds from these metadata, and formed the UI composition knowledge (¡component types¿¡bounding boxes¿) for each UI.
@@
Figure 2. Overview of PrototypeAgent System: This figure illustrates the main process of PrototypeAgent in response to a designer’s input and UI wireframe. PrototypeAgent utilizes a multi-agent approach for the interactive generation of UI prototypes. It encompasses four specialized agents—Theme Design $A_{theme}$, Textual Content $A_{text}$, Image Content $A_{img}$, and Icon $A_{icon}$. The Theme Design Agent $A_{theme}$ acts as the central coordinator, steering the collaborative efforts of the three sub-agents. By leveraging a cache pool, PrototypeAgent adeptly integrates the contributions from each agent to ensure a cohesive alignment with the overall design context. This process not only generates accurate prototypes but also provides explainable intermediate results, enabling designers to conduct thorough reviews and make precise edits.
@@ 4.2.2. Icon Knowledge Base
As visual shortcuts, icons improve the user experience in UI design. They enable intuitive navigation and enhance the visual appeal. To enhance a great generation capability and enable designers to modify based on their needs, we collected an icon knowledge base in SVG format for Icon Agent. This choice not only ensures compatibility across modern browsers but also guarantees that icons maintain their clarity when scaled. The designer can also modify the colour, style, and shape of the icons based on their specific requirements.
@@
4.3. Theme Design Agent
The Theme Design Agent functions as the central supervisor in the UI design process, leveraging domain knowledge from the UI knowledge base (Section 4.2). Its primary role is to perform implicit intent clarification through design prompt augmentation, setting the overall style and orchestrating the generation of a global theme description. By ensuring design coherence through the corresponding theme image, this dual-modality approach guarantees a high-quality UI design with consistency across all elements. The agent also collaborates with designers to align and refine design descriptions, facilitating a seamless design process.
Moreover, the Theme Design Agent works in tandem with a cache pool, collaborating with other agents to refine and detail specific components of the UI. The operational phases of the Theme Design Agent comprise four key stages: (1) Knowledge Retrieval: Accessing and utilizing domain-specific information. (2) Theme Description Generation: Facilitating implicit intent clarification and crafting a comprehensive and cohesive theme description. (3) Theme Image Generation: Producing a visual representation of the theme. (4) Sub-agent Execution: Coordinating sub-agents to execute tasks, ensuring intent-design alignment throughout the generation of each component. These stages are elaborated upon in subsequent sections, offering a detailed insight into each phase of the theme design process.
@@ 4.3.1. Knowledge Retrieval
Based on the design prompt ($In_{p}$) and UI layout ($In_{l}$), which is provided in a bounding box format accompanied by component labels as an example depicted in Fig. 4(B), we want to retrieve the most relevant knowledge from our large knowledge base. To do so, we concatenate these two information together, and encode them into one latent vector $\textbf{Emb}(In)$ as the query vector, where $In=In_{p}+In_{l}$. We use the TEXT-EMBEDDING-ADA-002 embedding model (OpenAi, 2024a)). Similarly, we also embed each piece of UI knowledge related to UI Composition and Semantic Understanding into a latent vector ($\textbf{Emb}(kb_{j})$) using the same embedding model. After that, we compute the cosine distance between the query vector and each knowledge vector, and retrieve the top-k results to instruct our multi-agent system. We denote the retrieved knowledge as $refer_{i}$.
@@ 4.3.2. Theme Description Generation
To illustrate the mechanisms driving this process, we detail a structured expression: Given a design prompt $In_{p}$, a UI wireframe layout $In_{l}$, and the top $k$ retrieved knowledge items $\{\sum_{i=0}^{k}refer_{i}\}$ (where, $k=2$), these components are concatenated with the system’s theme description prompt $P_{theme}$ to formulate a comprehensive input for our agent. Formally, the amalgamated prompt $P$ is delineated as:
@@
where $\oplus$ symbolizes concatenation, and the summation symbol $\sum$ here specifically indicates the sequential concatenation of retrieved knowledge items. With this integrated input at its disposal, the Theme Design Agent commences the generation of the Theme Description.
Upon completing the theme description generation, the resultant theme description is denoted by $Res_{theme}$. We show an example in Fig. 4, where the design prompt $In_{p}$ is depicted in part (A), the wireframe layout $In_{l}$ in part (B), the top $k$ retrieved knowledge items $\{\sum_{i=0}^{k}refer_{i}\}$ in part (C), and the system’s theme description prompt $P_{theme}$ is presented in part (D).

### v3

@@ 3.1. Findings
Figure 2. Illustrative Examples of Formative Study and Key Findings
@@
Our formative interviews confirmed the limitations of current input and output modalities used by designers, highlighting the need to address design aspects that better support the generation process. Unlike commercial tools, our PrototypeFlow focuses on prompt enhancement, enabling editable checkpoints during generation, and providing editable theme generation with downstream control. Table 1 in Section 2.1 compares popular design and generative tools used in participant workflows, helping to illustrate our findings and the advantages of our proposed method. We have identified five key findings as follows:
F1. Need for Streamlined Design Workflow Supported by Trend- and Brand-Aligned Knowledge
Participants reported a key challenge in moving from design ideas and low-fidelity wireframes to high-fidelity designs. This current process is often fragmented, requiring multiple tools and manual effort to obtain creative vision while aligning with company standards. As shown in Fig. 2 (F1), designers typically search online for implementation ideas, use tools like Midjourney for inspiration, and then consult internal guidelines to ensure consistency. This disjointed workflow—balancing trendiness, compliance, and diversity—demands considerable manual coordination. As one participant noted, “I regularly discover designs or elements that match what I’m looking for. However, seamlessly integrating them into my design process remains difficult. AI-powered tools should make this easier.” Participants also emphasized the need for scalable, evolving tools that can support generated UI design that reflect current trends and organizational styles.
System Design Response to F1: Motivated by these findings, our approach directly addresses these challenges using r etrieval-augmented generation (RAG), a dynamic, training-free method that leverages the knowledge base during the generation process. This knowledge base can be continuously updated to reflect evolving design trends and company-specific requirements, enabling the generated UIs to be not only creative and visually on-trend, but also aligned with concrete design constraints.
F2. Need for More Input Control and Flexible Output Editability in Design Generation Process
Participants expressed a shared frustration with existing AI tools, particularly regarding limited input control and output editability. They emphasized that both aspects are essential for efficiently, effectively, and accurately transforming their design requirements into deployable outcomes.
@@
Input Control/Modality: Traditional design tools often require manual creation of designs, typically in SVG format, which is time-consuming. Emerging generative tools allow designers to input natural language descriptions or high- and mid-fidelity design images (e.g., screenshots). However, interviews with participants revealed that being able to precisely and controllably describe their design intent is crucial. As one designer noted, “Text-only or image-only input doesn’t fully capture what I’m imagining, especially regarding the functionality of each component, which limits their practical use.” For example, a text prompt such as “Design a sleek, modern investment app welcome screen” can lead to layouts that ignore layout requirements. Conversely, image-only input may yield visually similar designs but also fails to capture pixel-level layout or functional interactivity (see “Bad Generation 1” and “Bad Generation 2” in Fig. 2, Input). Participants highlighted that combining wireframes with text descriptions allows for greater control over functionality and layout, while preserving creativity and the automation benefits of AI-powered tools.
@@
Output Editability/Modality: Our study found that while SVG-based outputs enable property-level editing and are designer-friendly, outputs provided as code (such as those from Vercel’s V0) or as static images (e.g., from MidJourney) make it difficult for designers to refine their work—especially for those without programming experience (as illustrated in Fig. 2, Output). As one participant shared, “Getting a design from MidJourney or Vercel’s V0 is amazing, but the real work begins with the day of editing that follows.” This highlights the need for output formats that are inherently editable and support iterative refinement.
System Design Response to F2: Guided by these findings, our system supports multimodal input by allowing designers to combine wireframes with natural language prompts, enabling specification of layout at the pixel level, functional precision, and explicit communication of design intent to achieve finer-grained input control. For output, the system generates editable SVG prototypes, ensuring that designers can directly manipulate visual properties and efficiently iterate on their designs. These features ensure that the design process remains both accurate and adaptable, fulfilling the identified requirements for modern design interaction modalities.
F3. Need for Supporting Designers in Expressing Intent Through Prompts
Our study found that designers often begin with simple and high-level prompts, such as “Design a sleek, modern investment app welcome screen.” without specifying the detailed requirements needed for an effective design. These initial prompts typically lack information about theme colors, categories, or other visual details. Designers expressed that, due to unfamiliarity with prompt structures, uncertainty about what the AI requires, and a desire to save time when writing prompts, they frequently provide insufficient detail. Participants emphasized that AI should play an active role in helping clarify and refine these initial prompts. As shown in Fig. 2 (F3), the system could automatically generate suggested prompts for theme settings, colors, or categories, and guide designers to provide more comprehensive requirements.
@@
As one participant explained, “I believe AI understands the kind of instructions it needs better than I do. If it could provide feedback, I could express my design goals more quickly and more accurately.” This highlights the importance of collaborative intent clarification, allowing designers and AI to reach a shared understanding of design goals, which aligns with the concept of achieving common ground in communication theory (Clark, 1991).
System Design Response to F3: To address this, our system includes a prompt enhancement step by generating specific suggestions for theme settings, colors, and other design properties, which are features lacking in existing AI generation tools (see comparison in Table 1, Prompt Enhancement). The system helps designers quickly specify detailed requirements, making the design process both faster and more accurate.
F4. Need for Precise Control in Generation Processes
Existing studies (Dang et al., 2023) highlight that end-to-end generation tools lack transparency and explainability, making it difficult for designers to align outputs with their true intent. As illustrated in Fig. 2 (F4), designers often receive prototypes that do not fully match their envisioned results. Without insight into the AI’s reasoning or step-by-step process, they are left to “guess” what prompt modifications will produce the desired outcome. This trial-and-error approach wastes time and reduces the effectiveness of AI-powered design.
@@
As summarized in Table 1 under Editable Checkpoint in Generation, current tools do not explain how outputs are generated or allow users to edit the reasoning process. Designers in our study called for more transparent and human-centered mechanisms to clarify and refine generation steps. As one participant remarked, “If AI tools could reveal their thought processes and allow editing, it would greatly simplify tailoring the results to our needs.”
System Design Response to F4: To address this, our system introduces transparent, editable checkpoints throughout the generation process. Designers can review and refine intermediate results instead of relying on trial and error. By adopting natural language as a domain-specific language (DSL) for global and local specifications, our approach enables collaborative, precise control over both style and components, leading to more accurate outcomes.
F5. Need for Maintaining Thematic Consistency and Coherence Across Generated Components
Maintaining visual consistency, such as color, style, and layout, across all components of a prototype is critical for professional design quality, but it is often repetitive and time-consuming. As illustrated in Fig. 2 (F5), AI generation tools may produce inconsistent components within a project, resulting in visual incoherence. Because of AI model hallucinations or forgetting, complex or lengthy pages can contain components that deviate from the intended theme. For example, as shown in the figure, although the overall theme color is blue, a green component may be generated. Traditional design tools do not provide mechanisms for editable theme generation with downstream control (see Table 1, Editable Theme Generation with Downstream Control), which forces designers to manually adjust inconsistencies.
One participant noted, “Creating a unified look for a new app involves a lot of repetitive work.” This highlights the need for tools that can automatically maintain design consistency and reduce the manual effort required for editing downstream components.
System Design Response to F5: To address this challenge, our system employs an LLM-based controller for editable theme generation with downstream control, ensuring consistency in color, style, and layout across all generated components. The controller transparently presents the natural language descriptions that govern design coherence, enabling designers to review, interact with, and efficiently refine consistency throughout the prototype.
@@ 4.1. Overview
To improve the findings identified in Section 3 and actualize the design space outlined, we introduce PrototypeFlow: a system that harnesses multi-module collaboration for the explainable generation of UI prototypes. This system aims to enhance mutual disambiguation in Human-AI Collaboration for creating high-fidelity prototypes by offering a decoupled generation process.
While large language models excel in contextual understanding and response generation, UI generation poses unique challenges that require deep domain knowledge to ensure high-quality design. A simple end-to-end approach may fall short of capturing the complexities of UI design intent. Therefore, we adopt a divide-and-conquer, top-down strategy that decouples the generation process into sub-steps. This modular approach not only clarifies the generation process for both designers and the system, but also supports mutual disambiguation in Human-AI collaboration.
As depicted in Fig. 3, PrototypeFlow is centrally orchestrated by the Theme Design Module $M_{theme}$, which coordinates three specialized modules: the Text Content Module $M_{text}$, the Image Content Module $M_{img}$, and the Icon Module $M_{icon}$. The system is supported by two curated knowledge bases—one for UI layouts (22k annotated screenshots) and another for 900 diverse icons—which ground the generative process in real-world design practices.
The theme design module serves as a central supervisor—similar to a centralized controller—that takes the user prompt and layout as input, retrieves relevant knowledge items from the knowledge base, generates a global theme description and component-level specifications, and produces a theme image to guide the image module. It ensures visual and textual consistency and coherence across the entire design, then sequentially invokes the corresponding sub-modules to generate high-fidelity content for each UI element. Finally, it assembles all generated components into the complete prototype. There is no interaction between sub-modules; all coordination and execution are managed centrally by the theme design module.
This modular, centrally-coordinated methodology enables PrototypeFlow to produce explainable, high-fidelity prototypes and supports iterative Human-AI collaboration. Designers benefit from both automatic generation and the ability to refine outputs, aligning results closely with their design intent.
@@ 4.2. Knowledge Base Construction
Employing domain-specific knowledge can harness the creativity of Large Language Models (LLMs) while enhancing the quality of generation (Wang et al., 2023c; Lu et al., 2024). We collected two kinds of knowledge, namely UI knowledge (pairs of layout/theme descriptions with local component descriptions) and Icon Knowledge (Pairs of icon SVG code with semantic descriptions) for $M_{theme}$ and $M_{icon}$.
@@ 4.2.1. UI Knowledge Base
1) UI Composition Knowledge. Rico Dataset (Deka et al., 2017) is one of the most comprehensive open-source UI datasets available, which contains around 22k distinct UIs from over 9.7k Android apps across 27 categories. This dataset includes the UI screenshots and their view hierarchy information, which expose UI elements used, their attributes like text, bounds and class, and the composition of these UI elements. We extracted class and bounds from these metadata, and formed the UI composition knowledge (¡component types¿¡bounding boxes¿) for each UI.
@@
Figure 3. Overview of PrototypeFlow System: This figure illustrates the main process of PrototypeFlow in response to a designer’s input and UI wireframe. PrototypeFlow utilizes a multi-modal approach for the interactive generation of UI prototypes. It encompasses four specialized modules—Theme Design $M_{theme}$, Textual Content $M_{text}$, Image Content $M_{img}$, and Icon $M_{icon}$. The Theme Design Module $M_{theme}$ acts as the central coordinator, steering the collaborative efforts of the three sub-modules. By leveraging a cache pool, PrototypeFlow adeptly integrates the contributions from each module to ensure a cohesive alignment with the overall design context. This process not only generates accurate prototypes but also provides explainable intermediate results, enabling designers to conduct thorough reviews and make precise edits.
@@ 4.2.2. Icon Knowledge Base
As visual shortcuts, icons improve the user experience in UI design. They enable intuitive navigation and enhance the visual appeal. To enhance a great generation capability and enable designers to modify based on their needs, we collected an icon knowledge base in SVG format for Icon Module. This choice not only ensures compatibility across modern browsers but also guarantees that icons maintain their clarity when scaled. The designer can also modify the colour, style, and shape of the icons based on their specific requirements.
@@
4.3. Theme Design Module
The Theme Design Module functions as the central supervisor in the UI design process, leveraging domain knowledge from the UI knowledge base (Section 4.2). Its primary role is to perform implicit intent clarification through prompt augmentation, setting the overall style and orchestrating the generation of a global theme description. By ensuring design coherence through the corresponding theme image, this dual-modality approach guarantees a high-quality UI design with consistency across all elements. The module also collaborates with designers to align and refine design descriptions, facilitating a seamless design process.
Moreover, the Theme Design Module works in tandem with a cache pool, collaborating with other modules to refine and detail specific components of the UI. The operational phases of the Theme Design Module comprise four key stages: (1) Knowledge Retrieval: Accessing and utilizing domain-specific information. (2) Theme Description Generation: Facilitating implicit intent clarification and crafting a comprehensive and cohesive theme description. (3) Theme Image Generation: Producing a visual representation of the theme. (4) Sub-module Execution: Coordinating sub-modules to execute tasks, ensuring intent-design alignment throughout the generation of each component. These stages are elaborated upon in subsequent sections, offering a detailed insight into each phase of the theme design process.
@@ 4.3.1. Knowledge Retrieval
Based on the design prompt ($In_{p}$) and UI layout ($In_{l}$), which is provided in a bounding box format accompanied by component labels as an example depicted in Fig. 5(B), we want to retrieve the most relevant knowledge from our large knowledge base. To do so, we concatenate these two information together, and encode them into one latent vector $\textbf{Emb}(In)$ as the query vector, where $In=In_{p}+In_{l}$. We use the TEXT-EMBEDDING-ADA-002 embedding model (OpenAi, 2024a)). Similarly, we also embed each piece of UI knowledge related to UI Composition and Semantic Understanding into a latent vector ($\textbf{Emb}(kb_{j})$) using the same embedding model. After that, we compute the cosine distance between the query vector and each knowledge vector, and retrieve the top-k results to instruct our multi-module system. We denote the retrieved knowledge as $refer_{i}$.
@@ 4.3.2. Theme Description Generation
To illustrate the mechanisms driving this process, we detail a structured expression: Given a design prompt $In_{p}$, a UI wireframe layout $In_{l}$, and the top $k$ retrieved knowledge items $\{\sum_{i=0}^{k}refer_{i}\}$ (where, $k=2$), these components are concatenated with the system’s theme description prompt $P_{theme}$ to formulate a comprehensive input for our module. Formally, the amalgamated prompt $P$ is delineated as:
@@
where $\oplus$ symbolizes concatenation, and the summation symbol $\sum$ here specifically indicates the sequential concatenation of retrieved knowledge items. With this integrated input at its disposal, the Theme Design Module commences the generation of the Theme Description.
Upon completing the theme description generation, the resultant theme description is denoted by $Res_{theme}$. We show an example in Fig. 5, where the prompt $In_{p}$ is depicted in part (A), the wireframe layout $In_{l}$ in part (B), the top $k$ retrieved knowledge items $\{\sum_{i=0}^{k}refer_{i}\}$ in part (C), and the system’s theme description prompt $P_{theme}$ is presented in part (D).